Liquid nitrogen has a density of 0.808 g/mL and boils at 77 K. Researchers often purchase liquid nitrogen in insulated 185 L tanks. The liquid vaporizes quickly to gaseous nitrogen (which has a density of 1.15 g/mL at room temperature and atmospheric pressure) when the liquid is removed from the tank. Suppose that all 185 L of liquid nitrogen in a tank accidentally vaporized in a lab that measured 10.00 m times 10.00 m times 3.50 m.
Part A
What maximum fraction of the air in the room could be displaced by the gaseous nitrogen?

Answer: Vgas/Vroom
Answer #1

Volume of room V(room)= 10.00 x 10.00 x 3.50 = 350 m3

Volume of liquid nitrogen = 185 L = 185 x 103 mL

Density of liquid nitrogen = 0.808 g/mL

Mass of nitrogen = volume x density

= 185 x 103 x 0.808 = 149480 g

Density of gaseous nitrogen = 1.15 g/L

Volume of gaseous nitrogen V(gas) = mass/density

= 149480/1.15 = 1.30 x 105 L = 130 m3

Fraction of air displaced

= V(gas)/V(room)

= 130/350 = 0.371